# The $\mathbb{Z}_{2}$-valued spectral flow of a symmetric family of Toeplitz operators

### v1

@@ Abstract.
We consider families $A(t)$ of self-adjoint operators with symmetry that causes the spectral flow of the family to vanish. We study the secondary $\mathbb{Z}_{2}$-valued spectral flow of such families. We prove an analog of the Atiyah-Singer-Robbin-Salamon theorem, showing that this secondary spectral flow of $A(t)$ is equal to the secondary $\mathbb{Z}_{2}$-valued index of the suspension operator $\frac{d}{dt}+A(t)$.
@@ 1. Introduction
The first main result of our paper is the analog the of the Atiyah-Singer-Robbin-Salamon theorem for half-spectral flow:
@@
Let $T_{f_{t}}^{\pm}$ denote the restriction of $T_{f_{t}}$ to $\mathscr{H}^{\pm}$. Under suitable assumptions on $D$ and $f_{t}$, cf. Section 7, the operators $T_{f_{t}}$ are Fredholm. We assume that the family $f_{t}$ is odd symmetric. Then so is the family $T_{f_{t}}$.
@@
We call this result the generalized bulk-edge correspondence because if $M=\{z\in\mathbb{C}:\,|z|=1\}$ is the unit disc in $\mathbb{C}$ this is equivalent to the bulk-edge correspondence for the Graf-Porta model for 2D topological insulators with time-reversal symmetry, [GrafPorta13]. In Section 9, we use (1.3) to obtain a new proof of the bulk-edge correspondence in the Graf-Porta model.
@@ 2.2. Odd symmetric operators
Let $W\subset H$ be a dense subspace and let $D:H\to H$ be an unbounded closed linear operator on $H$ whose domain is $W$. We endow $W$ with the inner product
@@ 3.1. Some spaces of operators
Let $A:H\to H$ be a closed operator with dense domain $W$. As in Section 2.2, we define the scala product on $W$ by
@@
Robbin and Salamon include an extra catenation axiom in their theorem. But then, [RobbinSalamon95, Prop. 4.26] they show, that this axiom follows from the others 4 axioms and, hence, can be omitted. Since for our main object of study – the half-spectral flow, º the catenation axiom does not make sense, we prefer to formulate Theorem 3.2 without it.
@@ 3.2. The spectral flow
$\operatorname{sf}A\ =\ \sum_{t}\operatorname{sign}\lambda(t).$ (3.4)
@@
    In every connected component of $\mathscr{A}(\mathbb{R},W,H)$ the is an $A\in\mathscr{A}^{1}$ which has only simple crossings.
@@ 4. The half-spectral flow of a family of operators on real line
Theorem 4.1 states that several properties of half-spectral flow define it uniquely. This is an analog of Theorem 3.2. If $A$ is a $\tau$-symmetric path, then the operator $D_{A}$, cf. (3.5) is $\tau$-symmetric (in the sense of Section 2) with respect to some anti-unitary anti-involution $\tau$. We use the uniqueness of the spectral flow to show that $\operatorname{ind}_{\tau}D_{A}$ is equal to the half-spectral flow of $A$. This is an analog of the Robbin–Salamon Theorem 3.6 for our secondary invariants. This result should be compared to [BourneCareyLeschRennie22, §8], though we consider different paths of operators.
4.1. An anti-linear involution
@@ 4.5. The $\tau$–index of $D_{A}$
We note that a similar version of the Robbin-Salamon result is obtained in [BourneCareyLeschRennie22, §8] for paths of skew-adjoint Fredholm operators on a real vector space.
@@
5. Proof of $\mathbb{Z}_{2}$ valued version of the Robbin-Salamon theorem
@@
Suppose $A_{1}$ and $A_{2}$ lie in the same connected component of $\mathscr{A}_{\tau}(\mathbb{R},W,H)$ and let $A_{s}$ ($s\in[0,1]$) be a continuous path in $\mathscr{A}_{\tau}$ connecting between them such that all $A_{s}$ have only isolated crossings. We need to prove that $\operatorname{sf}_{\tau}(A_{s})$ is independent of $s$. If $t=0$ is not a crossing point for some $A_{s_{0}}$, then there exists $\epsilon>0$ such that 00 is not a crossing for all $A_{s}$, $s\in(s_{0}-\epsilon,s_{0}+\epsilon)$. Then the path
@@
Since $t=0$ is not a crossing for $\tilde{A}_{s}$ ($s\in(s_{0}-\epsilon,s_{0}+\epsilon)$, $\operatorname{sf}_{\tau}(\tilde{A}_{s})$ is independent of $s$ on this interval.
@@ 5.3. A constant path
$\mu_{\tau}(A)\ =\ \mu_{\tau}(\tilde{A}_{\leq 0})\ =\ \mu_{2}(A_{\leq 0})\ \equiv\ \sum_{t<0}\operatorname{rank}P(A,t)\ =\ \operatorname{sf}_{2}(A).$ (5.4)
@@ 9.4. The edge Hamiltonian
We view $S^{1}\simeq[-\pi,\pi]$ as and involutive manifold with involution $\theta^{S^{1}}:t\to-t$. Then $H(t)$ is $\tau$-invariant by (9.6).

### v3

@@ Abstract.
We consider families $A(t)$ of self-adjoint operators with symmetry that causes the spectral flow of the family to vanish. We study the secondary $\mathbb{Z}_{2}$-valued spectral flow of such families. We prove an analog of the Atiyah-Patodi-Singer-Robbin-Salamon theorem, showing that this secondary spectral flow of $A(t)$ is equal to the secondary $\mathbb{Z}_{2}$-valued index of the suspension operator $\frac{d}{dt}+A(t)$.
@@ 1. Introduction
The first main result of our paper is the analog of the Atiyah-Patodi-Singer-Robbin-Salamon theorem for half-spectral flow:
@@
Note that recently a different version of the Robin-Salamon-type theorem for the $\mathbb{Z}_{2}$-valued spectral flow was obtained in [BourneCareyLeschRennie22]. In this paper, the authors considered paths of skew-adjoint Fredholm operators on a real vector space. The main difference with our result is that in [BourneCareyLeschRennie22] every operator $A(t)$ possesses a symmetry (it is skew-adjoint), while we consider the case when the symmetry relates $A(t)$ with $A(-t)$. We discuss the relationship between our result and [BourneCareyLeschRennie22] in Section 4.6.
@@
Let $T_{f_{t}}^{\pm}$ denote the restriction of $T_{f_{t}}$ to $\mathcal{H}^{\pm}$. Under suitable assumptions on $D$ and $f_{t}$, cf. Section 7, the operators $T_{f_{t}}$ are Fredholm. We assume that the family $f_{t}$ is odd symmetric. Then so is the family $T_{f_{t}}$.
@@
(Note that $\operatorname{sf}_{\tau}T_{f_{t}}^{-}=0$ since =$\mathcal{H}^{-}=\{0\}$). We call this result the generalized bulk-edge correspondence because if $M=\{z\in\mathbb{C}:\,|z|\leq 1\}$ is the unit disc in $\mathbb{C}$ this is equivalent to the bulk-edge correspondence for the Graf-Porta model for 2D topological insulators with time-reversal symmetry, [GrafPorta13]. In Section 9, we use (1.3) to obtain a new proof of the bulk-edge correspondence in the Graf-Porta model.
@@ 2.2. Odd symmetric operators
Let $W\subset H$ be a dense subspace and let $D:H\to H$ be a closed unbounded linear operator on $H$ whose domain is $W$. We endow $W$ with the inner product
@@ 3.1. Some spaces of operators
Let $A:H\to H$ be a closed operator with dense domain $W$. As in Section 2.2, we define the scalar product on $W$ by
@@
Robbin and Salamon include an extra catenation axiom in their theorem. But then, [RobbinSalamon95, Prop. 4.26] they show, that this axiom follows from the others 4 axioms and, hence, can be omitted. Since for our main object of study – the half-spectral flow – the catenation axiom does not make sense, we prefer to formulate Theorem 3.2 without it.
@@ 3.2. The spectral flow
$\operatorname{sf}A\ =\ \sum_{t}\operatorname{sign}\dot{\lambda}(t).$ (3.4)
@@
    In every connected component of $\mathscr{A}(\mathbb{R},W,H)$ there is an $A\in\mathscr{A}^{1}$ which has only simple crossings.
@@ 3.4. A periodic family of operators
Equivalently,
$\operatorname{ind}\left(\,\frac{d}{dt}+A\,\right)\ =\ \operatorname{sf}\tilde{A}.$ (3.9)
@@ 4. The half-spectral flow of a family of operators on real line
Theorem 4.1 states that several properties of half-spectral flow define it uniquely. This is an analog of Theorem 3.2. If $A$ is a $\tau$-symmetric path, then the operator $D_{A}$, cf. (3.5) is $\tau$-symmetric (in the sense of Section 2) with respect to some anti-unitary anti-involution $\tau$. We use the uniqueness of the spectral flow to show that $\operatorname{ind}_{\tau}D_{A}$ is equal to the half-spectral flow of $A$. This is an analog of the Robbin–Salamon Theorem 3.6 for our secondary invariants. This result should be compared to [BourneCareyLeschRennie22, §8], though we consider different paths of operators. We discuss the relationship of our result with [BourneCareyLeschRennie22] in Section 4.6.
4.1. An anti-linear anti-involution
@@ 4.5. The $\tau$–index of $D_{A}$
4.6. Comparison with the spectral flow of a path of skew-adjoint operators
For $A(t)\in\mathscr{A}_{\tau}(\mathbb{R},W,H)$ suppose for simplicity that the operator $A(0)$ is invertible. Set
$A_{<0}(t):\ =\ \begin{cases}A(t)\quad&\text{for}\ \ t\leq 0;\\ A(0)\quad&\text{for}\ \ t>0,\end{cases}\qquad\qquad A_{>0}(t):\ =\ \begin{cases}A(t)\quad&\text{for}\ \ t\geq 0;\\ A(0)\quad&\text{for}\ \ t<0.\end{cases}$
Consider the “doubling” of the path $A$:
$\hat{A}(t)\ :=\ \begin{pmatrix}0&A_{<0}(t)\\ -\alpha\,A_{>0}(-t)\,\alpha^{-1}&0\end{pmatrix}.$
This is a path of skew adjoint operators. The $\mathbb{Z}_{2}$-valued spectral flow of $\hat{A}$ considered in [BourneCareyLeschRennie22] is exactly equal to $\operatorname{sf}_{\tau}A$. Theorem 8.1 of [BourneCareyLeschRennie22] expresses this spectral flow as a $\mathbb{Z}_{2}$-valued index of a certain “suspension” operator of the form similar to $\frac{d}{dt}+A$. Using the gluing formula for the $\tau$-index, [BrSaeediYan25], one can show that the latter index is equal to $\operatorname{ind}_{\tau}\big{(}\frac{d}{dt}+A(t)\big{)}$. Thus the half-spectral flow of $A$ is equal to the $\tau$-index of $\frac{d}{dt}+A$. Combining it with Theorem 8.1 of [BourneCareyLeschRennie22], we see that the $\mathbb{Z}_{2}$-valued indexes of $\frac{d}{dt}+A$ and of the operator in [BourneCareyLeschRennie22] coincide. This can also be shown by a direct, but non-trivial argument.
In conclusion, our Theorem 4.4 is closely related to Theorem 8.1 of [BourneCareyLeschRennie22], but the relationship is not very straightforward. This is why we preferred to give a direct proof of our result in the next section, by adopting the original argument of Robbin-Salamon [RobbinSalamon95] to our $\mathbb{Z}_{2}$-valued invariants.
5. Proof of $\mathbb{Z}_{2}$-valued version of the Robbin-Salamon theorem
@@
Suppose $A_{0}$ and $A_{1}$ lie in the same connected component of $\mathscr{A}_{\tau}(\mathbb{R},W,H)$ and let $A_{s}$ ($s\in[0,1]$) be a continuous path in $\mathscr{A}_{\tau}$ connecting between them such that all $A_{s}$ have only isolated crossings. We need to prove that $\operatorname{sf}_{\tau}(A_{s})$ is independent of $s$. If $t=0$ is not a crossing point for some $A_{s_{0}}$, then there exists $\epsilon>0$ such that 00 is not a crossing for all $A_{s}$, $s\in(s_{0}-\epsilon,s_{0}+\epsilon)$. Then the path
@@
Since $t=0$ is not a crossing for $\tilde{A}_{s}$ ($s\in(s_{0}-\epsilon,s_{0}+\epsilon)$, $\operatorname{sf}(\tilde{A}_{s})$ is independent of $s$ on this interval.
@@ 5.3. A constant path
$\mu_{\tau}(A)\ =\ \mu_{\tau}(\tilde{A}_{\leq 0})\ =\ \mu_{2}(A_{\leq 0})\ \equiv\ \sum_{t<0}\operatorname{rank}P(A,t)\ =\ \operatorname{sf}_{\tau}(A).$ (5.4)
@@ 9.4. The edge Hamiltonian
We view $S^{1}\simeq[-\pi,\pi]$ as and involutive manifold with involution $\theta^{S^{1}}:t\to-t$. The anti-involution $\theta^{\mathbb{C}^{k}}$ defines a structure of a quaternionic bundle on the trivial bundle $S^{1}\times\mathbb{C}^{k}$ over $S^{1}$. As usual, we define an anti-involution
$\tau:\,L^{2}\big{(}S^{1},l^{2}(\mathbb{Z},\mathbb{C}^{k})\big{)}\ \to\ L^{2}\big{(}S^{1},l^{2}(\mathbb{Z},\mathbb{C}^{k})\big{)},\qquad\tau:\,f(t)\ \mapsto\ \theta^{\mathbb{C}^{k}}f(-t).$
Then $H(t)$ is $\tau$-invariant by (9.6).
@@ 10.1. The spectral flow as an index
Conflict of interest
The authors have no relevant financial or non-financial interests to disclose.
Data Availability
Data sharing is not applicable to this article as no new data were created or analyzed in this study.